### v1

@@ 3.3 Simple Event-based SL System Calibration
Figure 3: The calibration reprojection error of the projector and camera. The horizontal and vertical axes are measured in pixels.
Previous work[30] converts the event to images with E2VID [34] and utilize calibration tool [27] for rough extrinsic parameters calibration. Alternatively, some[28] adopt a step-by-step strategy and separately calibrate the camera’s intrinsic parameters. However, these methods (i) necessitates dual calibration procedures and complex calibration process, rendering it cumbersome to use. And (ii) images generated using E2VID may struggle to maintain sub-pixel level authenticity, while the step-by-step strategy introduces significant artificial accuracy loss. While some works have achieved high-precision calibration [39], they are limited to laser-point projectors or event cameras with gray mode.
Therefore, based on Gabriel’s calibration tool[8], we propose a novel, sub-pixel level, simple, and versatile event-based SL system calibration scheme SEC, as shown in the Fig. 2. Initially, we affix an LCD screen presenting a flickering chessboard onto a planar surface and project another chessboard pattern onto the other side. Following the acquisition of multiple datasets from various perspectives, the data is directly binarized and subjected to calibration procedures using the standard calibration tool. Our approach facilitates simultaneous calibration of all parameters through straightforward operations, significantly mitigating the complexity associated with calibration processes. Furthermore, our method attains high calibration precision, demonstrating the capability to achieve sub-pixel level reprojection errors. As evidenced in Fig. 3, we achieve the average reprojection error of [0.59, 0.54] for the projector and [0.40, 0.40] for the event camera in the experiment with 23 sets of data collected from different poses. Moreover, our SEC is universally applicable to all event-based SL systems.
@@
$D_{pr}$leftarrow ($D_{p}$, $C$) ;
// Obtain GX-map according to Equ. 6
$GX()$leftarrow $D_{pr}$ ;
@@
             $D_{cr}$leftarrow ($D_{c}$, $C$)
             // Disparity query according to Equ. 7 (Step (d))
             $dis$leftarrow ($D_{cr}$,$GX()$)
@@ 3.4 Depth Estimation
Figure 4: The projection characteristics of DLP projectors and laser-point projectors. (a) and (c) show the timestamp and time gradient variations with event id (sorted by time). (b) and (d) present the scatter plot of the events during the projection. DLP employs multi-point scanning, while laser-point uses raster scanning with non-linear characteristics. Both include dark intervals between projections..
GX-map Disparity Query. Following the acquisition of depth-encoded images, point matching is essential to extract depth information. Due to the high time-consuming nature of epipolar line search, we propose the GX-map disparity query scheme tailored for Gray code. Upon conducting stereo rectication to obtain rectified depth-encoded images, we store the x coordinates $x_{pr}$ of the projector’s rectified depth-encoded image over y coordinates $y_{pr}$ and the encoding value $v_{pr}$ as GX-map, then directly query the corresponding disparity $d$ using the rectified camera’s depth-encoded image. This is formulated as:
$v_{pr}=RP(x_{pr},y_{pr})\rightarrow x_{pr}=GX(y_{pr},v_{pr})$ (6)
$d=x_{cr}-GX(y_{cr},v_{cr})$ (7)
where $x_{cr}$, $y_{cr}$ and $v_{cr}$ represent the coordinates and encoding value of camera’s rectified depth-encoded image, respectively. $RP()$ and $GX()$ represent function that queries based on projector’s rectified depth-encoded image and GX-map. Given that the rectified depth-encoded image of the projector remains unchanged under identical intrinsic and extrinsic parameters, the GX-map generation operation requires execution only once and can subsequently facilitate multiple queries. Through GX-map querying, we have attained a thousand-fold enhancement in computational speed compared to traditional epipolar line search on the same processor, making high-speed real-time depth estimation possible.
@@
Time-overlapping. For dynamic scenes, due to the fact that a set of Gray codes starting with any serial number can compose a depth-encoded image, therefore, similar to[41], we propose a Time-overlapping strategy to maximize data utilization, as illustrated in Fig. 6.
@@
For depth estimation using N-bit Gray code, N projection patterns is required to compute depth. After capturing the initial set of binarized Gray codes, each subsequent Gray code can be combined with the previous images to calculate a new depth-encoded image. This means each projection corresponds to a new depth estimation, further improving the speed of the depth estimation by N times for N-bit Gray code.
@@ 4.1 Settings
In this task the ground truth is difficult to obtain because it is hard to find a method can generate dense depth of natural scenes above millimeter level accuracy. Therefore, ESL[30] uses averaged MC3D as their ground truth, while X-maps[28] uses optimized ESL. Others[25, 23] choose qualitative comparison only. Although we also think the averaging operation can improve the accuracy of depth estimation, which means averaged MC3D may be the most accurate method among others, we think because they both using the same point scanning method and vary only in different data processing process, therefore the comparison between them may also reflect the similarity of their depth results. This means the accuracy may be over-evaluated by selecting the same type of method as the baseline. Therefore it is unfair and impossible for our SGE to surpass other point scanning methods if we choose a point scanning method as ground truth.
But because our SGE uses different types of data, comparing our method with point scanning methods can fairly reflect our absolute precision. Therefore, we do not set ground truth in our comparison but compare our method with each point scanning method to verify the absolute accuracy of our scheme.
Metrics. The root mean square error (RMSE) and the fill-rate (FR) are adopted as our evaluation metrics. To minimize the influence of data differences, like slight differences in edge positions, we only use solid regions in both methods to calculate RMSE. We define solid regions as where both methods give a reliable depth estimation, which means the difference between them is within the threshold. The FR measures the ratio of points below depth error threshold in SGE to measured points in compared method. The depth error threshold is 1% of the average scene depth.
@@ 4.2 Static Scenes
Due to the inability of the point scanning schemes to utilize DLP projectors, we adopt the laser-point based system (as shown in Fig. 7 (b)) to eliminate device differences. In our experiment, we captured multiple scenes with varying complexity, including highly detailed gypsum statues, surfaces with fewer details, and everyday scenes to show the versatility of our solution. To minimize variables, we keep the projection light intensity, ambient light intensity, camera threshold, and shooting position constant while capturing two types of data. During the experiment, we use mean filtering to reduce discontinuity. Since the projector’s 720 horizontal resolution cannot be fully utilized by Gray code encoding, we use only 512 columns for projection, which results in the missing background on the right side.
Figure 8: Qualitative results of different methods on static scenes with the laser-point based system. Our SGE not only shows the performance consistent with the the-state-of-art schemes, but also presents more solid results in concave regions and less noise in overexposed and underexposed locations.
Table 1: Quantitative results of the differences between our SGE method and other methods in static scenes. RMSE and FR are adopted as evaluation protocol.
Scene David Dragon Structure gslim boxes fan bottle wizard arena column
Mean Depth 450mm 410mm 530mm 360mm 510mm 530mm 320mm 400mm 430mm 320mm
@@
MC3D-$1s$ 0.93 1.09 0.99 0.95 0.97 1.58 0.95 1.43 0.98 1.76 0.99 1.64 0.98 0.73 0.97 1.01 0.95 1.09 0.97 0.76
ESL 0.95 1.00 0.99 0.76 0.94 2.0 0.96 1.17 0.98 1.46 0.99 1.28 0.99 0.62 0.98 0.80 0.96 1.04 0.98 0.64
ESL-opt 0.95 0.99 0.99 0.90 0.97 1.66 0.95 1.40 0.98 1.70 0.99 1.62 0.99 0.68 0.98 1.00 0.96 1.12 0.99 0.72
Xmap 0.94 1.22 0.99 0.81 0.92 2.33 0.94 1.11 0.97 1.43 0.97 1.56 0.98 0.66 0.97 0.80 0.96 1.19 0.98 0.65
Fig. 8 presents the qualitative depth estimation results. It can be observed that for point scanning methods, due to timestamp noise and laser point diffusion, there is more error present at the concave regions (e.g. Structure, Boxes) and overexposed and underexposed locations (e.g. Wizard, Arena, Column). This is more pronounced in the X-map scheme, which does not utilize interpolation or averaging methods, and slightly improved in the ESL scheme with post-processing and in the MC3D scheme by averaging. However, our SGE demonstrates strong resistance to timestamp noise and laser point diffusion and performs accurate depth estimation in any case.
The quantitative results are shown in Tab. 1. Our method achieves a RMSE of less than 2mm, which is less than 1/200 of the average depth, and a FR over 0.93 when compared to other state-of-the-art methods. This indicates that our method achieves high accuracy comparable to the state-of-the-art with the absolute precision at the millimeter level. During the comparison, considering the potential application of real-time depth estimation, we just applied simple and fast median filtering in the pre-processing and post-processing stages to optimize the results. It is worth mentioning that our method can be further improved if additional off-line optimization algorithms are adopted, like those used in ESL.
@@ 4.4 Dynamic Scenes
Because obtaining recognized high-precision depths of natural scenes at a $\mu s$ resolution is difficult, if not impossible, without employing methods that rely on high-speed cameras or other costly equipment. Therefore, we cannot prove the accuracy of SGE directly. But as the experiment shown in Sec. 4.3 proves the high accuracy of our SGE at high scanning speeds, it can be assured that the accuracy of our scheme in dynamic scenes is guaranteed. The only additional factor impacting the accuracy is the incorrect depth encoding caused by motion, as the noise around the edge shown in Fig. 10. This is a problem that every dynamic scene depth estimation method will encounter. The only solution is to enhance data acquisition, which is the greatest strength of our SGE.
Figure 10: Visual results of our SGE on dynamic scenes. Various dynamic scenes at diverse timings are presented, proving our capacity of depth estimation on dynamic scenes
5 Conclusion
@@
In this work, We propose a high-speed event-based structured light depth estimation scheme: SGE. For the first time we introduce Gray code into event-based SL systems for depth estimation with a readily available DLP projector, realizing higher encoding efficiency, timestamp noise immunity and higher depth estimation speed. In addition, due to the absence of a simple and reliable calibration scheme, we present a sub-pixel level, universally applicable calibration method SEC for event-based SL system. Furthermore, we introduce the GX-map disparity query scheme and the Time-overlapping strategy to enhance computational speed and data utilization. We also captured a dataset consists of static and dynamic scenes with DLP and laser-point based system. Corresponding calibration data is also provided. Experimental results show that our SGE achieves kilohertz-level data acquisition speed while maintaining comparable accuracy as the state-of-the-art point scanning methods. This proves the capability of our SGE to realize high-speed and high-accuracy depth estimation.
In the future, we insist that our SGE can still be promoted by using more advanced event cameras with shorter refractory period and projectors with higher scanning speed. In addition, we believe a calibration scheme with a higher field of view (FoV) utilization and the study in reducing decoding errors caused by motion through data processing can also enhance our method.

### v2

@@ 3.3 Simple Event-based SL System Calibration
Previous work[20] converts the event to images with E2VID [58] and utilizes calibration tool [27] for rough extrinsic parameters calibration. Alternatively, some[21] adopt a step-by-step strategy and separately calibrate the camera’s intrinsic parameters. However, these methods (i) necessitate dual calibration procedures and a complex calibration process, rendering them cumbersome to use. And (ii) images generated using E2VID may struggle to maintain sub-pixel level authenticity, while the step-by-step strategy introduces significant artificial accuracy loss. While some works have achieved high-precision calibration [28], they are limited to laser-point projectors or event cameras with gray mode.
Therefore, based on Gabriel’s calibration tool[59], we propose a novel, sub-pixel level, simple, and versatile event-based SL system calibration scheme SEC, as shown in the Fig. 2. Initially, we affix an LCD screen presenting a flickering chessboard onto a planar surface and project another chessboard pattern onto the other side. Following the acquisition of multiple datasets from various perspectives, the data is directly binarized and subjected to calibration procedures using the standard calibration tool. Our approach facilitates simultaneous calibration of all parameters through straightforward operations, significantly mitigating the complexity associated with calibration processes. Furthermore, our method attains high calibration precision, demonstrating the capability to achieve sub-pixel level reprojection errors. As shown in Fig. 3, we achieve the average reprojection error and rough distribution range of [0.59, 0.54], [-2, 2] for the projector and [0.40, 0.40], [-1.5,1.5] for the event camera in the experiment with 23 sets of data collected from different poses. Moreover, our SEC is universally applicable to all event-based SL systems.
Figure 3: The calibration reprojection error of the projector and camera. The horizontal and vertical axes are measured in pixels. The average reprojection error and rough distribution range of [0.59, 0.54], [-2, 2] for the projector and [0.40, 0.40], [-1.5,1.5] for the event camera.
1
2
@@
4 $D_{pr}$leftarrow ($D_{p}$, $C$) ;
// Obtain GX-map according to Equ. 7
5 $GX()$leftarrow $D_{pr}$ ;
6
@@ 10             $D_{c}$leftarrow $G$
11             $D_{cr}$leftarrow ($D_{c}$, $C$)
             // Disparity query according to Equ. 8 (Step (d))
12             $dis$leftarrow ($D_{cr}$,$GX()$)
@@ 3.4 Depth Estimation
Figure 4: The projection characteristics of DLP projectors and laser-point projectors. (a) and (c) show the timestamp and time gradient variations with event id (sorted by time). (b) and (d) present the scatter plot of the events during the projection. DLP employs multi-point scanning, while laser-point uses raster scanning with non-linear characteristics. Both include dark intervals between projections.
@@
Figure 6: Schematic diagram of GX-map generation. Transform the rectified depth-encoded image of the projector that queries encoding values based on Y and X coordinate values into the GX-map that queries X coordinate values based on Y coordinate and encoding values. For example, the orange region are changed from “1 = RP(3,2)” to “2 = GX(3,1)”
GX-map Disparity Query. Following the acquisition of depth-encoded images, point matching is essential to extract depth information. Due to the high time-consuming nature of epipolar line search, we propose the GX-map disparity query scheme tailored for Gray code. Upon conducting stereo rectification to obtain rectified depth-encoded images of the projector and camera, we first store the x coordinates $x_{pr}$ of the projector’s rectified depth-encoded image over y coordinates $y_{pr}$ and the encoding values $v_{pr}$ to generate GX-map. As illustrated in Fig. 6, the left side is the rectified depth-encoded image of the projector, the right side is the corresponding GX-map, and the orange area represents the same data before and after conversion. Then, due to the fact that the matched points pair in rectified depth-encoded images of the projector and camera have the same y coordinates and encoding values, we can directly query the corresponding disparity $d$ using the rectified camera’s depth-encoded image. These are formulated as:
$v_{pr}=RP(y_{pr},x_{pr})\rightarrow x_{pr}=GX(y_{pr},v_{pr}),$ (7)
$d=x_{cr}-GX(y_{cr},v_{cr}),$ (8)
where $x_{cr}$, $y_{cr}$ and $v_{cr}$ represent the coordinates and encoding value of the camera’s rectified depth-encoded image, respectively. $RP()$ and $GX()$ represent functions that query based on the projector’s rectified depth-encoded image and GX-map. Given that the rectified depth-encoded image of the projector remains unchanged under identical intrinsic and extrinsic parameters, the GX-map generation operation only needs to be executed once and can then facilitate multiple queries. Through GX-map querying, we have attained a 200 times enhancement in computational speed compared to traditional epipolar line search on the same processor, making high-speed real-time depth estimation possible.
Time-overlapping. For dynamic scenes, since a set of Gray codes starting with any serial number can compose a depth-encoded image, therefore, similar to[60], we propose a Time-overlapping strategy to maximize data utilization, as illustrated in Fig. 7.
@@
For depth estimation using N-bit Gray code, N projection patterns are required to compute depth. After capturing the initial set of binarized Gray codes, each subsequent Gray code can be combined with the previous images to calculate a new depth-encoded image. This means each projection corresponds to a new depth estimation, further improving the speed of the depth estimation by N times for N-bit Gray code.
@@ 4.1 Settings
Accuracy assessment. In this task we want to show that our SGE achieves comparable accuracy to SOTA methods, therefore we compare our method with each of them to show the difference between them to verify the relative accuracy of our method instead of choosing one of SOTA methods as ground truth (GT). An available GT is difficult to obtain because it is hard to find a method that can generate dense depth of natural scenes above millimeter level accuracy, which is much more exact than all SOTA methods. Therefore previous works use similar different “reference” GT[20, 21], use different precision metrics[25, 26] or just qualitative comparison[19, 23]. Because our SGE and SOTA methods boast similar accuracy level, choosing one of them as GT may more exhibit the disparities in data types (point scanning vs. full-frame) and processing methods (time stamp matching vs. encoding value matching) rather than accuracy. Therefore we just compare the similarity between our scheme and SOTA methods, which is sufficient to demonstrate our comparable accuracy.
Metrics. The root mean square error (RMSE) and the fill-rate (FR) are adopted as our evaluation metrics. To minimize the influence of data differences, like slight differences in edge positions, we only use solid regions in both methods to calculate RMSE. We define solid regions as where both methods give a reliable depth estimation, which means the difference between them is within the threshold. The FR measures the ratio of points below the depth error threshold in SGE to measured points in the compared method. The depth error threshold is 1% of the average scene depth.
@@ 4.2 Static Scenes
Figure 9: Qualitative results of different methods on static scenes with the laser-point based system. Our SGE not only shows performance consistent with the the-state-of-art schemes but also presents more solid results in concave regions and less noise in overexposed and underexposed locations.
In the experiment, we captured multiple scenes with varying complexity, including highly detailed gypsum statues, surfaces with fewer details, and everyday scenes to show the versatility of our solution. To minimize variables, we adopt the laser-point based system (as shown in Fig. 8 (b)) to eliminate device differences due to the inability of the point scanning schemes to utilize DLP projectors. We also keep the projection light intensity, ambient light intensity, camera threshold, and shooting position constant while capturing two types of data. During the experiment, we employ the close operation as pre-processing in our SGE to fill in holes. Since the projector’s 720 horizontal resolution cannot be fully utilized by Gray code encoding, we use only 512 columns for projection, which results in the missing background on the right side in our SGE.
Fig. 9 presents the qualitative depth estimation results. During this experiment, we only use mean filtering to reduce discontinuity in all methods as post-processing. It can be observed that for point scanning methods, due to timestamp noise and laser point diffusion, there is more error present at the concave regions, e.g. Structure and Boxes, and overexposed and underexposed locations, e.g. Wizard, Arena, and Column. This is improved in the MC3D scheme by averaging. However, our SGE demonstrates strong resistance to timestamp noise and laser point diffusion and performs accurate depth estimation in any case.
Table 1: Quantitative results of the difference between our SGE and other methods in static scenes. RMSE and FR are adopted as evaluation protocols. Our method achieves an RMSE of less than 2.34mm, which is less than 1/200 of the average depth, and an FR over 0.91 compared to other methods.
Scene Bottle Dragon Structure Gslim Boxes Fan David Wizard Arena Column
Mean Depth 320mm 410mm 530mm 360mm 510mm 530mm 450mm 400mm 430mm 320mm
@@
MC3D-$1s$ 0.99 0.70 0.98 1.00 0.96 1.68 0.95 1.44 0.97 1.84 0.99 1.68 0.94 1.09 0.98 1.00 0.96 1.06 0.98 0.63
ESL 0.99 0.62 0.99 0.87 0.91 2.34 0.94 1.11 0.96 1.52 0.97 1.58 0.95 1.20 0.98 0.78 0.96 1.15 0.99 0.53
ESL-opt 0.99 0.62 0.98 0.95 0.94 1.94 0.94 1.20 0.95 1.76 0.97 1.63 0.95 1.10 0.98 0.97 0.96 1.10 0.99 0.55
Xmap 0.99 0.62 0.99 0.90 0.91 2.34 0.94 1.12 0.96 1.52 0.97 1.61 0.95 1.22 0.98 0.79 0.96 1.16 0.99 0.54
The quantitative results are shown in Tab. 1. Our method achieves an RMSE of less than 2.34mm, which is less than 1/200 of the average depth, and an FR over 0.91 compared to other state-of-the-art methods. This indicates that our method achieves high accuracy comparable to the state-of-the-art with precision at the millimeter level. During the comparison, considering the data gap and the fairness of the comparison, we utilize simple median filtering in the post-processing stage to optimize the results for all the methods. It is worth mentioning that our method can be further improved if additional off-line optimization algorithms are adopted at the cost of longer processing time, like those used in ESL.
@@ 4.4 Dynamic Scenes
Figure 11: Visual results of point scanning methods compared with our SGE on dynamic scenes. Various dynamic scenes at diverse timings are presented, proving our capacity of depth estimation on dynamic scenes. Given the similar qualitative performance across all point scanning methods, we use the X-map method as a representative. The areas marked by red boxes highlight the rolling shutter effect experienced with point scanning schemes, and those marked by brown boxes highlight the global shutter effect experienced with our SGE. Visualization 1-4 in certain dynamic scenes are also provided.
@@
We can observe that point-scanning approaches yield good results on low-speed dynamic scenes. However, for high-speed dynamic scenes, although they still produce relatively sharp edges, they suffer from the rolling shutter effect. In the sequence Tape, it is easy to observe that a three-dimensional rotating tape becomes twisted; Each blade of a rotating fan has a different size (e.g. Fan); Spheres become ellipsoids and the shape of the balls change during the falling and rebounding process in the sequence bounce. The fundamental cause of this effect is that the point-by-point scanning speed of point scanning schemes is lower than the speed of the object. Thus, increasing the scanning speed is the fundamental way to reduce the rolling shutter effect. However, point scanning schemes suffer from timestamp noise and low encoding efficiency, which hinders them from higher-speed depth data acquisition.
Although when the speed of dynamic scenes exceeds its capacity, our SGE is also affected by the "global shutter effect", which causes blur and noise around the edges of scenes. For example, as shown in Fig. 11, there is blur around the fan blade (e.g. Impact) and noise at the edges of the balloon (e.g. Balloon) and balls (e.g. Bounce). However, since our approach offers higher encoding efficiency and immunity to timestamp noise, we can realize higher depth data acquisition speed, thereby mitigating the impact of motion. This enables us to perform depth estimation for more rapid scenes which is unachievable by other point scanning methods. Our SGE has achieved satisfactory results in various challenging high-speed dynamic scenarios, such as multi-object interaction (e.g. Bounce, Collision, and Impact), thin structures at different speeds (e.g. Fan), and three-dimensional rotation (e.g. Tape). This proves that our SGE is capable of high-speed depth estimation, not only in simple scenes but also in various motion modes and complex scenes.
@@
Since it is difficult to obtain recognized high-precision depths of dynamic scenes at $\mu s$ temporal resolution without relying on high-speed cameras or other costly equipment, we are unable to directly prove the accuracy of SGE in such scenarios. But as the experiment shown in Sec. 4.3 proves the high accuracy of our SGE at high scanning speeds, it can be assured that the accuracy of our scheme in dynamic scenes is guaranteed. The only additional factor impacting the accuracy is the incorrect depth encoding caused by motion, as the noise and blur around the edge shown in Fig. 11. This is a problem that every dynamic scene depth estimation method will encounter. The fundamental solution is to enhance data acquisition, which is the greatest strength of our SGE.
4.5 Runtime evaluation for real-time processing
Table 2: Disparity calculation time (in seconds) on static scenes using our GX-map and traditional epipolar line search.
Scene David Dragon Structure gslim boxes fan bottle wizard arena column average
GX-map 0.08 0.08 0.08 0.08 0.08 0.09 0.08 0.09 0.08 0.09 0.08
Search 18.85 17.95 22.64 21.67 20.38 17.57 17.78 19.95 19.97 19.87 19.66 (246$\times$)
Table 3: The average total time required for a complete one-time depth estimation in dynamic scenes using both CPU and GPU (measured in seconds) of our SGE. Note that the maximum speed of other point scanning schemes is constrained by the point scanning speed of 60 Hz.
Scene Tape Balloon Collapse Fan_low Fan_mid Fan_high Bounce Collision Impact average
CPU 0.024 0.024 0.018 0.027 0.025 0.024 0.024 0.024 0.024 0.024 (42 Hz)
GPU 0.005 0.005 0.004 0.005 0.005 0.005 0.005 0.005 0.006 0.005 (200 Hz)
Additionally, we evaluate the computational cost of our SGE to illustrate the ability of our method to realize dense and high-precision depth estimation in real time. We use Python 3.9 on 12th Gen Intel i7-12700F CPU and RTX 3060 GPU to perform these experiments.
Tab. 2 shows the increase in speed using our GX-map disparity query. We evaluate the disparity calculation speed on static scenes using our GX-map and traditional disparity searching in CPU, the latter is the same algorithm ESL used. Two schemes use the same data as input and do not apply any algorithm optimization. It can be seen that our GX-map achieves around 250 times improvement in computational speed.
Tab. 3 shows the average total time of once depth map calculation. After simple algorithm optimization, our SGE realizes 200 Hz in GPU and 42 Hz in CPU for real-time depth estimation of 720 * 1280 resolution. It is worth mentioning that the X-map method, which has a similar computational complexity as ours, can also realize real-time processing, but the data acquisition speed of 60 Hz makes it impossible for them to perform faster real-time depth computing.
5 Conclusion, Limitation and Future Work
@@
In this work, we propose a high-speed event-based structured light depth estimation scheme: SGE. For the first time we introduce Gray code into event-based SL systems for depth estimation with a readily available DLP projector, realizing high depth acquisition speed while maintaining high accuracy. In addition, due to the absence of a simple and reliable calibration scheme, we present a sub-pixel level, universally applicable calibration method SEC for event-based SL systems. Furthermore, we introduce the GX-map disparity query scheme and the Time-overlapping strategy to enhance computational speed and data utilization. Experimental results show that our SGE achieves over kilohertz-level data acquisition speed and around 200 Hz real-time processing speed while maintaining comparable accuracy as the state-of-the-art point scanning methods. This proves the capability of our SGE to realize high-speed, high-accuracy, and real-time depth estimation.
Our method is primarily limited by the combined influence of the event camera sensor and illumination conditions, including ambient light and projected light intensity. While using a high projection light intensity and a high event camera threshold can almost entirely mitigate the effects of ambient light, depth-related uneven light intensity distribution, and motion events, this drastically slows the response speed of event cameras and lengthens the refractory period. Therefore, with current devices, we need to comprehensively adjust the ambient light intensity and projection light intensity at a lower event camera threshold to achieve the highest recovery accuracy. However, inappropriate light-intensity conditions may only cause a slight decrease in accuracy. Secondly, too fast motion in dynamic leads to depth estimation errors, and these errors will vary when updating the Gray code with different sequences using the Time-overlapping strategy. This can be mitigated by speeding up depth data collection using more advanced hardware. Finally, we believe that deeper optimization of the overall algorithm, the utilization of more advanced computing devices, and the substitution of more efficient programming languages can further enhance our real-time computing speed.
[…]